Journalist Report – February 1th

Crew 274 Journalist Report 2-1-2023

Sol: 3

Author: Tony DiBernardo, Crew Journalist

Title: Holocene and Harvest

The members of Crew 274 woke up this morning to the song “Holocene” by Bon Iver, sent to them by Dean Welsch, the Head of the School of STEM at APUS. His message outlined how appropriate the song is for our mission. “Holocene” is the name of our most recent geologic era of Earth’s history, where humans realize they are part of something bigger than themselves. What better than this to inspire us to continue our mission in preparing the future generations that will found a new settlement on Mars.

EVA Activity

After doing two EVAs on each of the first two sols of the mission, the crew took a break from EVAs today to focus on preparing tomorrow’s EVA to deliver the cache a 2-hour hike away for HSO Pender’s experiment.


Greenhab Officer Hines conducted daily research data shortly after conducting GreenHab operations at 11:00 am MST. Similar to the day prior, both microgreen and LDEF seed sets are retaining both temperatures and moisture while sustaining overnight temperatures. Additionally, the first signs of germination have been noted in both the cress and red cabbage microgreens, which will continue to be monitored closely over the upcoming days alongside the progress of the LDEF seed sets. To the


Chief Astronomer Pena and Heliophysicist Loy captured 70,000 images of heliophysic phenomena. These photos and videos supported critical space weather data and generated distinct, measurable visuals of the Sun’s outermost layer, the chromosphere, and solar dark spots, heliophysical granules, and solar prominences. This data was then analyzed, staked, measured, and configured through 3 astrophysics software applications. The resulting images will be used to help the astronomy crew in their scientific assessments of solar cycles and its implication to space weather.

Engineering and Hab Design

Hab Specialist O’Hara worked with Crew Engineer Lojek to troubleshoot and return the Musk Solar Observatory to nominal working order after experiencing a limit switch malfunction. O’Hara also completed an in-depth review of the Repair and Assembly Module (RAM) and the GreenHab, gathering photo documentation, and dimensional, concept of operations, traffic flow, systems and layout notes.

Astronomy Report – February 1th

Astronomy Report

Name: Salina Peña Noah Loy

Crew: 274

Date: 02/01/2023


Robotic Telescope Requested (Choose one) MDRS-14 MDRS-WF

Objects Viewed: Star/230201/variable star

Problems Encountered: None


Solar Features Observed: Sun/230201/sunspots

Sun/230201/solar prominence

Problems Encountered: The shutter issue from yesterday (01/31/2023) was resolved today and fully functioning now.

GreenHab Report – February 1th

GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 76.3 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 16.91

Daily water usage for research and/or other purposes: 2.11 gallons

Water in Blue Tank: 155.88 gallons

Time(s) of watering for crops: 11:00 am

Changes to crops: N/A

Narrative: Continued to monitor and water GreenHab crops during mid-morning operations. Additional watering and attention were given to the tomato plants with additional trimming of debris and dead leaves from each plant to support continued growth and daily harvest.

Harvest: <1 gram of dill, 1 gram of thyme, 13 grams of carrots, 37 grams of microgreens, 55 grams of beans, 41 grams of tomato, 5 grams of oregano, 22 grams of spinach, and 13 grams of basil.

Support/supplies needed: N/A

Operations Report – February 1th

Crew 274 Operations Report 2/1/2023

SOL: 3

Name of person filing report: Alexis Lojek

Non-nominal systems: Oven light; toilet (now nominal)

Notes on non-nominal systems: Oven light still needs a new bulb. We previously reported that the specifications of the oven light bulb type were emailed to mission support by crew 272 engineer. The toilet required draining into the septic tank this morning, so that was completed. It is now nominal.


Spirit rover used: No

Hours: 207.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: 219.5

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: 254.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

General notes on rovers: None of the four rovers were used on EVA. They are all charging from last night.

Summary of Hab operations: Habitat is in a nominal condition.

WATER USE: 23 gallons

Water (static tank): 393 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Radios were charged after mission completion yesterday and taken off the chargers after the red charging lights turned off. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants this morning at 1100-1130 am.

WATER USE: 17 gallons used.

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: 186 grams

Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is fully setup continuing to be monitored.

Dual split: On.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The Musk Solar Observatory shutter doors were returned to nominal operating state by the engineers. They checked the voltage on the battery and manually actuated the switches, which fixed the issue.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: The propane alarm went through its full startup cycle again when power swapped over to the generator in the evening. This seems to happen whenever power is swapped over from one power source to another.

Sol Summary Report – February 1th

Crew 274 Sol Summary Report 02-01-2023

Sol: 3

Summary Title: Solar Phenomena

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: The crew did not perform any EVAs today. The day was used to plan out the logistics cache study, opening the solar telescope, harvesting green beans, tomatoes, and herbs. The crew also used the day to clean the mud from the lower deck, vacuum the crew quarters, and clean the bathroom. The solar study was a massive success. Crew Astronomers Loy and Pena captured 70,000 images. The images showed sunspots, prominences, and granules.

Look Ahead Plan: The crew looks forward to possible cache EVA deployment and new solar images from the astronomy crew.

Anomalies in work: None

Weather: Sunny, -2C High, -9C Low

Crew Physical Status: Satisfactory

EVA: None

Reports to be filed: Sol Summary, Operations, EVA Request, HSO, GHO Report, and Astronomy Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site